OpenStudy (kaylala):

Trigonometric Equation 2cosA - 3tanA = 0

hartnn (hartnn):

write tan as sin/cos

OpenStudy (kaylala):

Okay. I will try that

hartnn (hartnn):

then it will become similar to your previous question

OpenStudy (kaylala):

But there's sin this time How's that? @hartnn

hartnn (hartnn):

previous to previous, similar to the one that had cos^2 A and sin A

OpenStudy (kaylala):

But this has no square root @hartnn

hartnn (hartnn):

http://openstudy.com/users/kaylala#/updates/52ec9b98e4b0cdd2a983b48d same formula you need to use

OpenStudy (kaylala):

I got nothing. REally sorry @hartnn

hartnn (hartnn):

2c -3t =0 2c -3s/c =0 2c^2 -3s =0 use c^2+s^2 =1

OpenStudy (kaylala):

2c^2 -3s =0 How did you get that?

hartnn (hartnn):

multiplying c on both sides

OpenStudy (kaylala):

Okay

OpenStudy (kaylala):

Will it be x^2+3x+2?

hartnn (hartnn):

x^2+3x-2=0 isn't it ?

OpenStudy (kaylala):

Oh houre right

OpenStudy (kaylala):

*you're

OpenStudy (kaylala):

What's next? I really dont know

OpenStudy (kaylala):

@hartnn ^

hartnn (hartnn):

solve the quadratic eq.

OpenStudy (kaylala):

Right. Just a moment

OpenStudy (kaylala):

X= (-3+-√(17))/2 Is that correct?

hartnn (hartnn):

the quadratic eq. should be 2x^2+3x-2=0

OpenStudy (kaylala):

Oh my bad

OpenStudy (kaylala):

Let me try again

OpenStudy (kaylala):

A=30,150 Is that it?

OpenStudy (kaylala):

@hartnn

OpenStudy (kaylala):

Sin A =-2 Has no solution Right?

hartnn (hartnn):

yes, correct 30,150 also correct :)

OpenStudy (kaylala):

Great help. Thanks again @hartnn